# Generating and Understanding Tests via Path-Aware Symbolic Execution with LLMs

### v1

@@ III-B Path Aware Test Generation with Validation
After extracting path variants, PALM organizes them into a symbolic execution tree, where each leaf node corresponds to a complete path variant, and internal nodes represent shared prefixes between paths. PALM performs test generation by traversing this tree in a top-down fashion. Algorithm 3 formalizes this process. At each node, in Line 7-12, PALM sends the corresponding path variant directly to the LLM for test generation, such as the path variant shown in Fig. 4(c). For each path, PALM allows up to five trials to generate a assertion-satisfying test and backtracks if the path is deemed infeasible. This symbolic execution tree not only facilitates test generation, but also provides the foundation for checking whether a specific path is covered and for visualizing test coverage at the path level.
@@
1:$root$: The root node of the program’s symbolic execution tree
2:$T$: A test suite aligned with program paths
3:function GenerateTests($node$)
4:    $p\leftarrow$ the path condition from $root$ to $node$
5:    if $node$ is a leaf then
6:       Attempt to generate a test for path $p$ (up to 5 trials)
7:       Return $\{t\}$triangleright A valid test if exists
8:    end if
9:    if $node$ is divergent then
10:       Run test generation for path $p$
11:       if unsatisfiable or generation fails after 5 trials then
12:          return $\emptyset$
13:       end if
@@
15:    return $\bigcup\textsc{GenerateTests}(c)$,   $\forall c\in child(node)$
16:end function
TABLE I: PALM’s prompt template for path-aware test generation. The prompt is designed to guide the test generation task to a specific program path by encoding its constraints as a sequence of assertTrue and assertFalse statements. PALM attempts test generation iteratively, and upon failure, it provides error feedback to the LLM by returning the first divergent branch between the generated test and the target path.
@@
When constructing the prompt for test generation, PALM retains the full implementation of class fields and non-symbolic methods from the input program. This provides essential context for the LLM to reason about (1) input-output behavior and (2) internal state transitions within non-symbolic functions. PALM excludes the implementation of external or third-party library functions not present in the input code, as modern LLMs are already well-trained on common APIs. Including such definitions would be redundant and could dilute prompt clarity by increasing its length without contributing additional semantic value. Table I shows PALM’s prompt template during test generation.
This ability to leverage LLMs eliminates the need to symbolically model used library functions—a major pain point of traditional symbolic execution engines, such as Symbolic Pathfinder [2], which require developers to either manually encode each function’s semantics as symbolic constraints or treat them as uninterpreted.
Path exercisability validation with runtime test execution
When the LLM returns a test, PALM executes it against the corresponding program variant to verify whether it indeed exercises the intended path. If the test fails, PALM identifies the first violated assertion and feeds it back to the LLM for test refinement.
For instance, given the path in Fig. 4(d), the LLM may initially generate the test is_palindrome("ab"). Execution shows that the assertion at 6 fails because the second iteration should evaluate the condition to true. PALM returns this failure assertTrue(text.charAt(i) != text.charAt(len-i-1)) to the LLM, which then produces a corrected test is_palindrome("abca") that successfully exercises the intended path.
III-C Implementation
We implemented the path extraction component based on JavaParser [8], and developed the frontend using Vue.
IV Evaluation
@@
  • RQ1: Does PALM’s path-aware program variant generation improves path coverage?
  • RQ2: Does PALM’s path exercisability validation and error-feedback guided test generation improves path coverage?
  • RQ3: Does PALM overcome the limitation of traditional symbolic executors (e.g. Symbolic Pathfinder) on handling external functions without symbolic modeling?
@@
We used 124 benchmark programs from HumanEval-Java, a Java benchmark for evaluating automated program repair tools [7]. Its programs contain various external method calls, data structure manipulations, and complex control flow.
@@
In our evaluation of PALM, we set the loop unrolling bound to 2 and the recursion depth bound to 2. We treat the method whose name matches the target class as both the symbolic function and the entry point. All other methods—whether defined within the program or invoked from external APIs—are treated as non-symbolic and are not subject to path enumeration. During test generation on the symbolic tree, we enumerate up to 50 distinct paths and terminate further test generation once this limit is reached.
IV-A Path-aware test generation
1public static boolean any_int(double x, double y, double z) {
2 if ((int) x == x && (int) y == y && (int) z == z) {
3 if (x + y == z || x + z == y || y + z == x)
4 return true;
5 } // else-branch
6 return false;
7}
Figure 5: Code snippet from any_int. The condition (int)x == x checks whether x is an integer (i.e., has no fractional part). The highlighted else-branch corresponds to inputs where at least one of x, y, or z is not an integer. LLM-generated tests (GPT-4o-mini) fail to cover this branch, whereas PALM covers with any_int(3.0,1.1,2.0).
TABLE II: Comparison of average test coverage between PALM and LLM under different LLM backends. Each configuration was executed five times to reduce variance. Reported cost represents the total for all 124 benchmarks.
Tool Path Coverage Branch Coverage Line Coverage Cost (\$)
@@ IV-A Path-aware test generation
Table II illustrates PALM’s code coverage against LLM. PALM with GPT-4o-mini achieves a 35.0% higher path coverage, 5.9% branch coverage, 4.0 % line coverage than direct LLM generation with GPT-4o-mini. It is worth noting that PALM with GPT-4o-mini achieves higher path coverage than direct LLM generation with a stronger model GPT-o3-mini. PALM with GPT-o3-mini achieves a 24.2% higher path coverage, 1.7% lower branch coverage, 0.4% lower line coverage than direct LLM generation with GPT-o3-mini.
GPT-4o-mini achieves 8.8%, 1.1%, and 0.4% higher path, branch, and line coverage than PALM, respectively. This result is due to the path limit we set in PALM to avoid path explosion. By default, PALM only generates tests for the first 50 paths and skips the rest. The visualization interface of PALM helps users identify which parts of the program do not need path enumeration, giving users more control over what to test.
IV-B Test validation and refinement
11112222333344445555600600600600700700700700Path Cov@K11112222333344445555190190190190195195195195200200200200Branch Cov@K11112222333344445555940940940940950950950950Line Cov@K
GPT-4o-miniGPT-o3-miniGPT-o4-mini
Figure 6: Test coverage progress of PALM with $k$ rounds of trial for each program path using three LLM backends
@@ IV-B Test validation and refinement
IV-C Comparison with Symbolic PathFinder
TABLE III: Java code patterns that lack symbolic support in Symbolic PathFinder (SPF). For these cases, SPF either terminates prematurely or falls back to concrete execution without complete path modeling.
@@ IV-C Comparison with Symbolic PathFinder
Generic Types HashSet<?>.add(), HashMap<?,?>.put() 32 (22.4%) Hard to model e.g. Object.equals, Object.compareTo for generic objects
@@
We compared PALM against Symbolic PathFinder. SPF, as typical for many symbolic execution engines, requires symbolic modeling of used API functions or considers them as uninterpreted.
@@ 3 String last = " ";
Figure 7: SPF lacks sufficient symbolic modeling for the two highlighted API calls: split and toLowerCase. The converted SMT path constraints are thus imprecise. In contrast, PALM utilizes LLM’s capability to interpret the semantics of commonly used APIs.
Fig. 7 shows an example, where SPF fails to model the path constraints due to two unsupported API calls and to model a symbolic heap object. Symbolic modeling is particularly challenging in this case because SPF must model the symbolic heap object from split function, which is a String[] array—an object type that SPF currently does not support symbolically. On the other hand, falling back to concrete execution on split() loses the dependency between txt and parts, resulting in an incomplete path modeling.
Table III summarizes the list of library functions that have not been modeled by SPF. As a result, SPF fails to model even a single complete path in 49 programs (34.3%). In contrast, PALM does not lower path constraints into SMT. PALM adopts symbolic execution-style path enumeration and utilizes LLM to infer constraints such as parts=txt.split(" "), bypassing the need for explicit symbolic modeling.
IV-D User study
TABLE IV: Demographic breakdown of user study participants
Participant Type Count
CS PhD Student 6
CS Master Student 5
CS Undergraduate Student 1
Total 12
To understand how PALM assists users in generating path-aware tests, we conducted a controlled user study comparing PALM against a baseline LLM-based approach. Our goal was to evaluate whether PALM can help users better understand and identify redundant or missing test cases in terms of different path profiles and produce concrete tests aligned with specific paths. We reported users’ task accuracy, completion time, and confidence when using each tool.
Our user study follows a counterbalanced, within-subject design, not a between-subject design. While between-subject user studies require a large number of participants to account for variations among individual participants, within-subject user studies are not affected by variability, as each participant tries out two different tools in a different order. The order of tool usage and task assignment in within-subject studies is randomized and counterbalanced.
@@ IV-D User study
We recruited 12 participants, whose backgrounds is detailed in Table IV. All of them were familiar with Java and had experience with using LLMs.
TABLE V: Programs and questions used in user study.
@@
Q2 Same Path Test From a list of candidate tests, identify the one that exercises the same execution path as a given reference test.
Q3 Targeted Test From a list of candidate tests, select a test that would cover a particular program path described in natural language.
@@
To evaluate how each tool supports path-aware test generation, participants were asked to generate tests for two subject programs using either PALM a pure LLM chatbox. As a part of test generation task, they completed three targeted comprehension questions designed to assess their understanding of (1) the total number of feasible paths (path enumeration), (2) the equivalence between tests in terms of path exercising behavior, and (3) the ability to match a test to a natural-language path description. Table V summarizes the subject programs and associated questions.
@@
We adopted a 2$\times$2 crossover design [9], where each participant used both tools (PALM and baseline) on two different programs in varied orders, evenly distributed to mitigate order bias and learning effects.
@@
TABLE VI: Post-study questionnaire on tool confidence
@@
We used a post-task questionnaire to obtain feedback. Table VI lists the three questions to assess participants’ perceived confidence in completing test generation tasks using PALM or LLMs. All three questions adopt a 5-point Likert scale, where 1 indicates ”Not confident” and 5 indicates ”Very confident.”
TABLE VII: Comparison of participant task accuracy and task completion time using different tools
@@
00555510101010w/ PALMP1: Path Coverage Confidence00555510101010w/ LLM00555510101010w/ PALMP2: Redundancy Avoidance Confidence00555510101010w/ LLM00555510101010w/ PALMP3: Targeted Test Generation ConfidenceNot confident (1)Slightly confident (2)Moderately confident (3)Fairly confident (4)Very confident (5)00555510101010w/ LLM
Figure 8: Post-study questionnaire results: distribution of confidence scores (x-axis) and corresponding participant counts (y-axis) for using PALM or LLM.
@@
As shown in Table VII, participants using PALM achieved higher accuracy across all three questions: 0.92 on Q1 and 1.0 on both Q2 and Q3. The only incorrect response occurred on Q1, which assessed path enumeration, where one participant answered 6 instead of the correct 7. This participant was observed to correctly navigate the symbolic tree, which visualizes all feasible paths and allows users to inspect which ones are covered or missing, but still miscounted the total number. This suggests a simple counting error rather than confusion about the number or structure of feasible paths.
In contrast, participants using the LLM chatbox found it harder to identify all feasible paths and determine which tests exercised which paths. On Q1, four participants answered that there were only four feasible paths in CuteArray, underestimating the correct count of seven. We observed that the LLMs they used often missed edge test cases where the loop executes zero or one time—scenarios that are critical for full path coverage. On Q3, which required selecting a test input that matches a natural-language path description, one participant chose the wrong test from a list of similar candidates. She copied the description directly into the prompt and accepted the first response without further inspection. The returned test diverged from the intended path at the inner condition of the second loop iteration due to a misunderstanding of the data dependency. In comparison, participants using PALM could verify which tests exercised a specific path through the visualized execution trace.
Completion time. In terms of time, participants using PALM spent slightly more on average (5.43 minutes vs. 5.17 minutes for LLM). However, 1.68 minutes of this was spent on automated test generation, requiring no user effort. The remaining time was used for reviewing results and answering questions. In contrast, participants using the LLM chatbox had to craft prompts, interpret outputs, and manually verify whether the tests aligned with the path mentioned in the task.
Participant behavior under the LLM condition varied. Some quickly accepted the first model response without checking, resulting in shorter times but more errors, while some spent more time validating results.
@@
We observed that PALM helped participants better understand which paths were covered or missing (P1), enabling them to augment their test suites to cover missing cases. It also improved their understanding of how each test input exercises a different execution path (P2, P3), allowing them to identify which paths are covered or not. This was particularly useful in reducing redundant testing effort.
V Related Work
V-A Symbolic Execution
Symbolic execution is a popular program analysis technique introduced to test whether certain properties can be violated by a software component [10, 11, 12, 13]. A key challenge in symbolic execution is to model system and environment behaviors under the limitation of constraint solving [14]. Earlier works such as DART [15] and CUTE [16] adopt concolic execution by using fully concrete arguments for those external functions. This leads to insufficient path exploration and incomplete symbolic modeling of path constraints within concretely executed functions. Another line of works create symbolic abstractions for such external APIs. KLEE [1] has a basic symbolic file system model. AEG [17] symbolically models library calls related to the file system, network sockets, and environment variables. CLOUD9 [18] provides additional symbolic abstractions for POSIX libraries.
In contrast, Palm does not rely on explicit symbolic abstractions. Instead, it preserves path semantics at the higher Java language level, leveraging LLM’s test generation capabilities towards specific programs with inserted assertions. This approach eliminates the need to define complex symbolic abstractions in SMT format and is suitable for handling external API functions that are closed-source.
V-B LLM Test Input Generation

### v2

@@ 3.2. Path Aware Test Generation with Validation
After extracting path variants, PALM organizes them into a symbolic execution tree, where each leaf node corresponds to a complete path variant, and internal nodes represent shared prefixes between paths. PALM performs test generation by traversing this tree in a top-down fashion. Algorithm 3 formalizes this process. Line 7-12 send the corresponding path variant to the LLM for test generation. For each path, PALM allows up to five trials to generate an assertion-satisfying test and backtracks if the path is deemed infeasible. This symbolic execution tree not only facilitates test generation, but also provides the foundation for checking whether a specific path is covered and for visualizing test coverage at the path level.
@@
0:rootroot: the root node of the program’s symbolic execution tree
0:$T$: a test suite aligned with program paths
1:Procedure GenerateTests(nodenode)
[…]
3:if nodenode is a leaf then
4:  $t\leftarrow$ attempt to generate a test for path $p$ (up to 5 trials)
5:  if $t$ is not found then
6:   return $\emptyset$
7:  end if
8:  return $\{t\}$ // Return a valid test if exists
9:end if
10:if nodenode is divergent then
11:  Run test generation for path $p$
12:  if unsatisfiable or generation fails after 5 trials then
13:   return $\emptyset$
@@
15:end if
16:return $\bigcup\texttt{GenerateTests}(c)$, $\forall c\in child(node)$
17:End Procedure
Table 1. PALM prompt template. Path constraints are encoded as assertTrue/assertFalse statements, and failures provide assertion-level feedback for iterative refinement.
A table showing the PALM prompt template with four labeled sections: (1) Task Description instructing the LLM to generate Java tests that satisfy all assertTrue and assertFalse statements without printing or extra logic; (2) Input program listing placeholders for [imports], [other functions], and the [path condition]; (3) Generation History recording the current round number, previous [test code] and any [failed assertion] used for iterative refinement; and (4) Output Format specifying the expected test output. Placeholders shown in the table (for example [focal method]) are meant to be substituted when constructing a path-specific prompt.
@@
When constructing the prompt for test generation, PALM retains the full implementation of class fields and non-symbolic methods from the input program. This provides essential context for the LLM to reason about (1) input-output behavior and (2) internal state transitions within non-symbolic functions. PALM excludes the implementation of external or third-party library functions not present in the input code, as modern LLMs are already well-trained on common APIs. Including such definitions would be redundant and could dilute prompt clarity by increasing its length without contributing additional semantic value. Table 1 shows PALM’s prompt template during test generation.
By leveraging LLMs for test generation, PALM eliminates the need to symbolically model used library functions, a major pain point of traditional symbolic execution engines such as Symbolic Pathfinder (Păsăreanu and Rungta, 2010). In such engines, developers must either manually encode each function’s semantics as symbolic constraints or treat them as uninterpreted.
Path exercisability validation via runtime execution
When the LLM produces a test, PALM runs it on the corresponding program variant to check whether it exercises the target path. If it fails, PALM reports the first violated assertion to the LLM for refinement. For example, for the path in Fig. 4(d), the LLM may generate is_palindrome("ab"). The run violates the assertion at 6 (the second iteration should evaluate the condition to true). PALM feeds back assertTrue(text.charAt(i) != text.charAt(len-i-1)), and the LLM corrects the test to is_palindrome("abca"), which exercises the intended path.
4. Evaluation
@@
  • RQ1: Does PALM’s path-aware program variant generation improve path coverage?
  • RQ2: Does PALM’s path exercisability validation and error-feedback guided test generation improve path coverage?
  • RQ3: Does PALM overcome the limitation of traditional symbolic executors (Symbolic Pathfinder) on handling external functions without symbolic modeling?
@@
We used 124 benchmark programs from HumanEval-Java, a Java benchmark for evaluating automated program repair tools (Jiang et al., 2023). Its programs contain various external method calls, data structure manipulations, and complex control flow.
Implementation
We implemented the path extraction component based on JavaParser (JavaParser Team, 2025), and developed the frontend using Vue.
@@
In our evaluation, we set the loop-unrolling and recursion-depth bounds to 2. We treat the method whose name matches the target class as the symbolic entry, and treat all other methods (including program-defined helpers and external API calls) as non-symbolic and excluded from path enumeration. During test generation, we enumerate up to 50 distinct paths per program (including infeasible paths), capping exploration to mitigate path explosion and bound analysis cost. We chose 50 as a practical budget that keeps the total cost bounded while covering the majority of feasible paths in HumanEval-Java. Increasing this limit leads to diminishing returns in path coverage relative to the additional LLM cost. Each experiment is run 5 times.
4.1. Path-aware test generation
Table 2. Comparison of average test coverage between PALM and LLM under different LLM backends. Each configuration was executed five times to reduce variance. Reported cost represents the total for all 124 benchmarks.
A table comparing average test coverage and total cost for PALM and a direct LLM baseline across three LLM backends: GPT-4o-mini, GPT-o3-mini, and GPT-o4-mini. For each backend, the table reports average path coverage, branch coverage, and line coverage, along with the total dollar cost for running all 124 benchmarks, with each configuration repeated five times.
Tool Path Cov Branch Cov Line Cov Cost (\$)
@@ 4.1. Path-aware test generation
1public static boolean any_int(double x, double y, double z) { 2 if ((int) x == x && (int) y == y && (int) z == z) { 3 if (x + y == z || x + z == y || y + z == x) 4 return true; 5 } // else-branch 6 return false; 7}
Figure 5. Code snippet from any_int. The condition (int)x == x checks whether x is an integer (i.e., has no fractional part). The highlighted else-branch corresponds to inputs where at least one of x, y, or z is not an integer. LLM-generated tests (GPT-4o-mini) fail to cover this branch, whereas PALM covers with any_int(3.0,1.1,2.0).
A formatted Java code listing for the method any_int(double x, double y, double z). The code first checks whether x, y, and z have no fractional part using comparisons such as (int)x == x, and then tests whether any pairwise sum equals the third value. An else-branch is highlighted to indicate the case where at least one input is not an integer, after which the method returns false.
Table 2 illustrates PALM’s code coverage against LLM. PALM with GPT-4o-mini achieves a 35.0% higher path coverage, 5.9% branch coverage, 4.0 % line coverage than direct LLM generation with GPT-4o-mini. It is worth noting that PALM with GPT-4o-mini achieves higher path coverage than direct LLM generation with a stronger model GPT-o3-mini. PALM with GPT-o3-mini achieves a 24.2% higher path coverage, 1.7% lower branch coverage, 0.4% lower line coverage than direct LLM generation with GPT-o3-mini.
We also observe that, as the LLM backend becomes stronger, direct generation improves and can even outperform PALM under our current configuration: with GPT-o4-mini, direct generation achieves 8.8%, 1.1%, and 0.4% higher path, branch, and line coverage than PALM, respectively. This is partly because PALM caps exploration at 50 enumerated paths per program to mitigate path explosion. Nevertheless, PALM remains complementary via path-level guidance and test-path visualization for diagnosing missing or redundant coverage.
4.2. Test validation and refinement
Figure 6. Test coverage progress of PALM with $k$ rounds of trial for each program path using three LLM backends
A line chart showing test coverage progress of PALM as the number of trial rounds \$k\$ increases for each program path. The chart compares three LLM backends, with separate curves illustrating how coverage improves over successive rounds for each backend.
@@ 4.2. Test validation and refinement
4.3. Comparison with Symbolic PathFinder
Table 3. Java code patterns that lack symbolic support in Symbolic PathFinder (SPF). For these cases, SPF either terminates prematurely or falls back to concrete execution without complete path modeling.
A table summarizing Java code patterns that lack symbolic support in Symbolic PathFinder (SPF). The rows list four categories—String operations, type conversion, generic types, and other library calls—each with example code structures, the number and percentage of affected benchmark programs, and a brief explanation of why SPF fails (for example, missing symbolic models for String APIs such as split(), lack of SMT support for string-to-double conversion, difficulty modeling generic-object methods such as Object.equals and Object.compareTo, and fallback to concrete execution for unsupported library calls).
@@ 4.3. Comparison with Symbolic PathFinder
Generic Types HashSet<?>.add(), HashMap<?,?>.put() 32 (22.4%) Hard to model method calls for generic objects, e.g. Object.equals and Object.compareTo
@@
We compared PALM against Symbolic PathFinder (SPF). As is typical for many symbolic execution engines, SPF requires symbolic modeling of used API functions or considers them as uninterpreted.
@@ 3 String last = " ";
Figure 7. SPF lacks sufficient symbolic modeling for the two highlighted API calls: split and toLowerCase. The converted SMT path constraints are thus imprecise. In contrast, PALM utilizes LLM’s capability to interpret the semantics of commonly used APIs.
A Java code listing for a method check_if_last_char_is_a_letter(String txt) that splits the input string on spaces, selects the last token, normalizes it, and checks whether the first character is a letter from a to z. Two API calls are highlighted: txt.split(" ") and Character.toLowerCase(last.charAt(0)).
Fig. 7 shows an example, where SPF fails to model the path constraints due to two unsupported API calls and to model a symbolic heap object. Symbolic modeling is particularly challenging here because SPF would need to represent the heap object returned by split(), namely a String[] array, a type that SPF does not currently support symbolically. Alternatively, falling back to concrete execution on split() loses the dependency between txt and parts, resulting in an incomplete path modeling.
Table 3 summarizes the list of library functions that have not been modeled by SPF. As a result, SPF fails to model even a single complete path in 49 programs (34.3%). In contrast, PALM does not lower path constraints into SMT. PALM adopts symbolic execution-style path enumeration and utilizes LLM to infer constraints such as parts=txt.split(" "), bypassing the need for explicit symbolic modeling.
4.4. User study
To understand how PALM assists users in generating path-aware tests, we conducted a controlled user study comparing PALM with a baseline LLM-based approach. We evaluate whether PALM helps users better understand path coverage, identify redundant or missing tests, and produce concrete tests aligned with specific execution paths. We report task accuracy, completion time, and self-reported confidence for each tool. Our user study follows a counterbalanced, within-subject design. Each participant uses both tools in different orders, with tool usage and task assignment randomized and counterbalanced to mitigate learning and order effects.
@@ 4.4. User study
We recruited 12 participants: 6 CS PhD students, 5 CS master’s students, and 1 CS undergraduate student. All participants were familiar with Java and had prior experience using LLMs.
Table 4. Programs used in user study.
A table listing the programs used in the user study, with columns for subject name, lines of code (LOC), number of paths, and a brief description. The table includes two subjects: CuteArray (12 LOC, 7 paths), which counts divisible-by-7 triplets in a loop while modifying array elements, and Triangle (19 LOC, 12 paths), which sorts three side lengths and checks whether they form a valid triangle.
@@
Table 5. Questions used in user study.
A table listing the questions used in the user study, with columns for question ID, question title, and a brief description. The table includes three questions: Q1 on path enumeration within a fixed loop bound, Q2 on identifying a test that exercises the same execution path as a reference test, and Q3 on selecting a test that covers a specific path described in natural language.
@@
Q2 Same Path Test From a list of candidate tests, identify the one that exercises the same path as a given reference test.
Q3 Targeted Test From a list of candidate tests, select a test that would cover a particular path described in natural language.
@@
To evaluate how each tool supports path-aware test generation, participants were asked to generate tests for two subject programs using either PALM or a pure LLM chatbox. As a part of test generation task, they completed three targeted comprehension questions designed to assess their understanding of (1) the total number of feasible paths, (2) the equivalence between tests in terms of path exercising behavior, and (3) the ability to match a test to a natural-language path description. Table 4 and Table 5 summarize the subject programs and associated questions.
@@
We adopted a $\times2$ crossover design (Deljouyi et al., 2024), where each participant used both tools (PALM and baseline) on two different programs in varied orders, evenly distributed to mitigate order bias and learning effects.
@@
Table 6. Post-study questionnaire on tool confidence
A table showing the post-study questionnaire items measuring tool confidence, with columns for item ID, question title, and description. The items ask participants how confident they are that the tool achieves comprehensive path coverage (P1), avoids generating redundant tests for the same path (P2), and generates a test that exercises a specified target path (P3).
@@
We used a post-task questionnaire to obtain feedback. Table 6 lists the three questions to assess participants’ perceived confidence in completing test generation tasks using PALM or LLMs. All three questions adopt a 5-point Likert scale, where 1 indicates ”Not confident” and 5 indicates ”Very confident.”
Table 7. Comparison of participant task accuracy and task completion time using different tools
A table comparing user-study results for two tools, PALM and a direct LLM baseline. The table reports participant task accuracy for three questions (Q1, Q2, and Q3) and the average task completion time in minutes. PALM shows accuracies of 0.92 for Q1 and 1.0 for Q2 and Q3 with an average time of 5.43 minutes, while the LLM baseline shows accuracies of 0.58 for Q1, 1.0 for Q2, and 0.92 for Q3 with an average time of 5.17 minutes.
@@
As shown in Table 7, participants using PALM achieved higher accuracy across all three questions: 0.92 on Q1 and 1.0 on both Q2 and Q3. The only incorrect response occurred on Q1, which assessed path enumeration, where one participant answered 6 instead of the correct 7. This participant was observed to correctly navigate the symbolic tree, which visualizes all feasible paths and allows users to inspect which ones are covered or missing, but still miscounted the total number.
Figure 8. Post-study questionnaire results: distribution of confidence scores (x-axis) and corresponding participant counts (y-axis) for using PALM or LLM.
A stacked set of bar charts showing post-study confidence ratings on a 1-5 Likert scale for three questions (P1 path coverage confidence, P2 redundancy avoidance confidence, and P3 targeted test generation confidence). For each question, two horizontal panels compare responses with PALM versus with a direct LLM baseline, with the y-axis showing participant counts (0-12). Most PALM responses concentrate at high confidence levels (4-5), while LLM responses are more spread across levels 1-5.
In contrast, participants using the LLM chatbox struggled to enumerate feasible paths and match tests to paths. On Q1, four participants underestimated the number of feasible paths in CuteArray (4 vs. 7), often missing cases where the loop executes 0 or 1 iteration. On Q3, one participant selected an incorrect candidate after directly prompting the description and accepting the first response, leading to a mismatch at an inner condition due to overlooked data dependency. In comparison, PALM users could verify test-path alignment via the visualized execution trace.
Completion time. Participants using PALM and participants using the LLM chatbox completed tasks in a similar time range (5.43 vs. 5.17 minutes on average).
@@
We observed that PALM helped participants better understand which paths were covered or missing (P1), enabling them to augment their test suites to cover missing cases. It also improved their understanding of how each test input exercises a different execution path (P2, P3), allowing them to identify which paths are covered. This was particularly useful in reducing redundant testing effort.
5. Threats to Validity
Internal validity.
Our comparison with direct LLM prompting is not budget-equivalent. PALM enumerates paths and performs path-specific generation and validation, which typically requires longer prompts and higher LLM usage than one-shot prompting. Thus, PALM’s coverage gains may partly reflect more generation and validation opportunities. We do not include an ablation that equalizes LLM budgets, so results reflect PALM under its intended workflow rather than a cost-controlled comparison.
For Symbolic PathFinder (SPF), we align key search settings such as the same loop-unrolling bounds. However, SPF outputs satisfying assignments rather than executable unit tests and does not synthesize JUnit-style tests from these assignments, especially for arrays or heap-allocated objects. Therefore, we do not report SPF path coverage on HumanEval-Java; instead, we report how many programs yield at least one SMT-formatted path constraint as a coarse applicability indicator.
External validity.
We evaluate PALM on HumanEval-Java, whose programs are relatively small and primarily use Java standard-library APIs. This setting may not fully represent real-world projects with larger codebases, richer third-party dependencies, and more complex behaviors arising from I/O, concurrency, reflection, and framework-driven callbacks. Moreover, our implementation targets Java, and the results may not generalize to other languages. Finally, HumanEval benchmarks may be subject to data contamination in LLM training corpora, which could inflate the absolute performance of LLM-based approaches.
Construct validity.
We primarily use branch-based path coverage as the quantitative metric. This metric does not capture data-flow behaviors that do not manifest as distinct branch outcomes and may under-represent the strengths of traditional symbolic execution in generating boundary values within the same branch structure. Our user study complements coverage by evaluating users’ ability to reason about path-level behaviors using PALM’s interface. However, the study tasks focus on matching tests to paths and identifying redundant or missing tests, which may differ from end-to-end usage of test-generation tools in real projects. We also do not comprehensively measure other test-quality dimensions such as readability, naturalness, and bug-finding capability.
Conclusion validity.
Our user study includes 12 participants, which limits statistical power and the strength of quantitative claims. Since LLM-based generation is stochastic, single-run results can be sensitive to sampling variance; we reduce this risk by repeating runs and summarizing aggregate results, but residual variance may still exist.
6. Related Work
Symbolic execution is a widely used program analysis technique for systematically reasoning about program behaviors and generating concrete inputs that witness property violations (King, 1975; Boyer et al., 1975; King, 1976; Howden, 1977). A key challenge is modeling system and environment interactions under the constraints of solver-based reasoning (Baldoni et al., 2018). DART (Godefroid et al., 2005) and CUTE (Sen et al., 2005) provide concolic execution to execute external functions concretely, which can under-approximate path constraints and limit exploration when important conditions reside inside those calls. To mitigate this, several systems introduce symbolic abstractions for common external APIs, e.g., KLEE’s symbolic file system (Cadar et al., 2008), AEG’s models for file systems, network sockets, and environment variables (Avgerinos et al., 2014), and Cloud9’s POSIX abstractions (Bucur et al., 2011).
In comparison, PALM avoids translating path constraints into SMT formulas and therefore does not require SMT-level symbolic abstractions for external functions. However, the current implementation enumerates paths primarily by explicit branch outcomes. In contrast, mature symbolic execution engines (Păsăreanu and Rungta, 2010) can additionally explore paths induced by runtime exceptions and implicit checks (e.g., nullness, bounds, divide-by-zero), as well as thread-scheduling nondeterminism in concurrent programs.